### v1

@@ Abstract
The interplay between topology and quantum criticality gives rise to the notion of symmetry-enriched criticality, which has attracted considerable attention in recent years. However, its non-Hermitian counterpart remains largely unexplored. In this Letter, we show how parity-time ($\mathcal{PT}$) symmetry enriches non-Hermitian critical points, giving rise to a topologically distinct non-unitary universality class. By analytically investigating non-Hermitian free fermion models with $\mathcal{PT}$ symmetry, we uncover a new class of conformally invariant non-unitary critical points that host robust topological edge modes. Remarkably, the associated topological degeneracy is surprisingly encoded in the purely imaginary part of the entanglement entropy scaling—a feature absent in Hermitian systems. The underlying mechanism for the emergence of edge states at non-Hermitian criticality is traced to a generalized mass inversion that is absent in Hermitian systems.
@@
Introduction.—Non-Hermitian (nH) quantum systems have recently attracted considerable attention due to their unique properties beyond Hermitian counterparts [1, 2, 3, 4, 5] and are closely related to diverse experimental platforms [6, 7, 8, 9, 10]. The characterization of quantum phases in nH systems is now well established [11, 12, 13, 14, 15, 16], revealing distinctive phenomena such as the non-Hermitian skin effect [17, 18, 19, 20, 21, 22] and non-Hermitian topological phases [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. In contrast, phase transitions in nH systems—typically described by non-unitary conformal field theories (CFTs)—remain less understood, although they have recently drawn growing interest from the perspective of quantum entanglement [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70].
@@
On a different front, the universality class of quantum phase transitions can be further enriched by global symmetries, giving rise to topologically distinct universality classes, now referred to as symmetry-enriched quantum criticality (SEQC) [71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88] or, more generally, gapless symmetry-protected topological (gSPT) states  [89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121]. The discovery of topological physics in quantum critical systems opens new avenues for classifying phase transitions within the same universality class, fundamentally enriching the textbook understanding of phase transitions and going beyond the traditional paradigms of statistical and condensed matter physics. Remarkably, the coexistence of symmetry-protected edge states with a gapless bulk gives rise to intriguing topological phenomena absent in gapped topological phases, including nontrivial conformal boundary conditions [76, 92], algebraically localized edge modes [71, 117], universal bulk-boundary correspondence [76, 102, 83], and intrinsically gapless topological phases [96, 115].
A fundamental open question is whether the notion of SEQC can be generalized to non-unitary criticality in non-Hermitian systems, and if so, what mechanisms govern such phenomena. Despite its importance, this issue has remained largely unexplored due to the intrinsic difficulty of systematically investigating quantum criticality in generic non-Hermitian many-body systems.
@@
To make progress in answering the above questions, we systematically explore $\mathcal{PT}$-symmetry–enriched non-unitary criticality in a family of one-dimensional nH free-fermion models, uncovering a new class of non-unitary critical points that host robust topological edge modes. Using the combination of exact solution and numerical simulations, we show that although both critical points in these models are described by non-unitary CFTs with central charge c=2c=-2, one of them is further enriched by $\mathcal{PT}$ symmetry, giving rise to topologically non-unitary criticality with stable edge states. Remarkably, the associated topological degeneracies are encoded in the purely imaginary part of the entanglement entropy scaling, a feature unique to non-Hermitian systems. Finally, we elucidate the emergence of topological edge states at this critical point by proposing an entirely new mechanism—generalized mass inversion—that is absent in Hermitian counterparts.
@@
which depends only on $v_{k}$. The vertical dashed line $w=v$ in Fig. 1(a) therefore marks a change in edge-mode count that is independent of $u$. Note that inside the $\mathcal{PT}$-broken wedge the dashed line is not a phase boundary, because the Berry (Zak) phase, which determines the phase structure, no longer corresponds to the winding number there [122].
@@
First we review the ordinary ($\alpha=1$) nH SSH model. At its trivial QCP, the correlation eigenvalues are real and occur in $(\nu,1{-}\nu)$ pairs, ensuring that the entanglement entropy remains real and scales with central charge c=2c=-2, consistent with a non-unitary CFT (see Fig. 2(a)).
In contrast, the topologically nontrivial QCP exhibits a complex conjugate pair of correlation eigenvalues, $\nu_{\pm}=1/2\pm i\mathcal{I}$, which lead to an imaginary contribution to the entanglement entropy. While previous studies [36] encountered ambiguities due to the multivalued complex logarithm, we restore the correct analytic structure by consistently shifting one logarithmic branch by $2\pi$. This yields an additional term $i\pi$ in the entropy, while preserving the scaling form with c=2c=-2 in the real part (see Fig. 2(b)). Unlike in Hermitian systems, the entropy here acquires a constant imaginary part; we defer its physical interpretation to a later section. Full details of the branch-cut prescription are provided in the Supplementary Material [122].
@@
Entanglement–edge correspondence and the imaginary entropy—At the critical points of the $\alpha$-nH SSH chain, we observe a correspondence between the physical and entanglement spectra. Under OBC, the system supports localized edge modes with purely imaginary energies $E=\pm iu$. Under PBC, the entanglement spectrum features complex-conjugate eigenvalue pairs of the form $\nu=1/2\pm i\mathcal{I}$, whose number matches that of the OBC edge modes [see Fig. 3(a,b)]. Using the standard correlation–entanglement mapping $\nu=(1+e^{\epsilon})^{-1}$, equivalently
@@
corresponding to a purely imaginary entanglement energy. This generalizes the Li–Haldane picture: although both energy are no longer real or degenerate, the one-to-one counting and the vanishing real parts ($\operatorname{Re}[E]=\operatorname{Re}[\epsilon]=0$) still match.
@@
A key consequence concerns the imaginary part of the entanglement entropy. Requiring the correct c=2c=-2 scaling at all critical points forces us to adopt the same branch-cut convention used in the previous section. With this convention, each complex-conjugate pair contributes $i\pi$ to $S_{A}$. Together with the entanglement spectrum–edge correspondence, at a QCP with winding number $\omega$,
@@
This identifies a subleading imaginary contribution $i\pi\omega$. Unlike in Hermitian SEQC, where topological features at criticality are reflected only in the entanglement spectrum [101, 83], the nH case renders them directly visible in the imaginary part of the entanglement entropy—a feature unique to nH systems.
It is then natural to extend the same branch choice into the adjacent SPT phase by continuity. With that extension, $\operatorname{Im}[S_{A}]$ takes the same constant value throughout each fixed-$\omega$ sector. Thus $\operatorname{Im}S_{A}$ serves as a topological index reflecting the winding number $\omega$ across the $\mathcal{PT}$-symmetric sector.
@@
Generalized mass inversion in $\mathcal{PT}$ symmetry-enriched non-Hermitian QCP—A distinctive feature of the $\alpha$-nH SSH model is the appearance of an intermediate $\mathcal{PT}$-broken symmetry-breaking phase absent in Hermitian free-fermion systems. This additional phase shifts the location of the critical point and is closely tied to the persistence of edge modes across the transition. In the Hermitian $\alpha\geq 2$-chain, topological edge states rely on kinetic inversion with long-range hopping to maintain edge states at criticality. By contrast, in the nH SSH model, the survival of edge modes follows from a simple generalization of mass inversion without requiring long-range hopping ($\alpha=1$).
@@
decouple at $E=\pm iu$, yielding localized edge solutions with decay length $\sim 1/m$ for $m>0$ even as the bulk gap $\sqrt{m^{2}-u^{2}}$ closes.
@@
More generally, this is due to when we have more than one mass operators $\{m_{a}\Gamma_{a}\}$ with $\Gamma_{a}$ being the gamma matrices in the Dirac-like Hamiltonian. Here some of the $m_{a}$ can be imaginary. Edge localization is set by the domain-wall mass—the component that flips across the boundary—so $\xi\sim 1/|m_{bdy}|$. The bulk gap, however, is controlled by the spectral mass $\Delta=\sqrt{\sum_{a}m_{a}^{2}}$. Thus these scales are able to decouple when we have multiple mass term. Including this scale decoupling together with ordinary mass inversion is what we refer to as generalized mass inversion.
This mechanism is absent in Hermitian systems, where the gap is always enhanced and cannot close without delocalizing the edge modes. This makes a free-fermion SEQC without long-range hopping possible only in the nH setting. For higher $\alpha$, a generalized mass inversion coexists with kinetic inversion, see the SM [122].
@@
The schematic in Fig. 4(b) depicts an interface geometry (left: trivial gapped phase; right: nH SSH at its topological QCP). Unlike the OBC case, the interface equations remain coupled and lack a simple closed-form solution; nevertheless, generalized mass inversion pins an edge state at the interface [122].
@@
Discussion and concluding remarks.—To summarize, we uncover a new class of non-unitary criticality enriched by $\mathcal{PT}$ symmetry, which hosts robust topological edge states. Specifically, we construct a broad family of one-dimensional $\mathcal{PT}$-symmetric nH free-fermion models with $\alpha$-range hopping, all exhibiting critical points described by non-unitary CFTs with central charge c=2c=-2, as confirmed numerically through the scaling of entanglement entropy and ground-state energy. More importantly, we unambiguously demonstrate that these nH critical points support nontrivial edge states protected by $\mathcal{PT}$ symmetry. The associated topological degeneracy is reflected in the bulk entanglement spectrum—representing a generalized Li-Haldane correspondence at nH criticality—and, remarkably, is also encoded in the imaginary part of the entanglement entropy, a feature unique to nH systems. Finally, we analytically show that the topological edge modes at these QCPs are stabilized by a new mechanism, generalized mass inversion—without requiring long-range hopping—highlighting a distinct protection mechanism from Hermitian counterparts.
Looking ahead, identifying symmetry-enriched nH criticalities in higher dimensions and in interacting many-body systems represents an interesting direction for future work. Regarding experimental realization, we note that nH free-fermion lattice models and entanglement-based observables can be implemented and probed using quantum and classical simulators, including ultracold atoms [123, 124], superconducting quantum processors [125, 126], and phononic systems [127]. Our work not only generalizes the notion of topology to the realm of non-unitary criticality but also contributes to a deeper understanding of critical phenomena in nH systems.

### v2

@@ Abstract
The interplay between topology and quantum criticality gives rise to the notion of symmetry-enriched criticality, which has attracted considerable attention in recent years. However, its non-Hermitian counterpart remains largely unexplored. In this Letter, we show how parity-time ($\mathcal{PT}$) symmetry enriches non-Hermitian critical points, giving rise to a topologically distinct non-unitary universality class. By analytically investigating non-Hermitian free fermion models with $\mathcal{PT}$ symmetry, we uncover a new class of conformally invariant non-unitary critical points that host robust topological edge modes. Remarkably, the associated topological degeneracies are surprisingly encoded in the purely imaginary part of the entanglement entropy scaling—a feature absent in Hermitian systems. The underlying mechanism for the emergence of edge states at non-Hermitian criticality is traced to a generalized mass inversion that is absent in Hermitian systems.
@@
Introduction.—Non-Hermitian (nH) quantum systems have recently attracted considerable attention due to their unique properties beyond Hermitian counterparts [1, 2, 3, 4, 5] and are closely related to diverse experimental platforms [6, 7, 8, 9, 10]. The characterization of quantum phases in nH systems is now well established [11, 12, 13, 14, 15, 16], revealing distinctive phenomena such as the non-Hermitian skin effect [17, 18, 19, 20, 21, 22] and non-Hermitian topological phases [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. In contrast, phase transitions in nH systems—typically described by non-unitary conformal field theories (CFTs)—remain less understood, although they have recently drawn growing interest from the perspective of quantum entanglement [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74].
@@
On a different front, the universality class of quantum phase transitions can be further enriched by global symmetries, giving rise to topologically distinct universality classes, now referred to as symmetry-enriched quantum criticality (SEQC) [75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92] or, more generally, gapless symmetry-protected topological states  [93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125]. The discovery of topological physics in quantum critical systems opens new avenues for classifying phase transitions within the same universality class, fundamentally enriching the textbook understanding of phase transitions and going beyond the traditional paradigms of statistical and condensed matter physics. Remarkably, the coexistence of symmetry-protected edge states with a gapless bulk gives rise to intriguing topological phenomena absent in gapped topological phases, including nontrivial conformal boundary conditions [80, 96], algebraically localized edge modes [75, 121], universal bulk-boundary correspondence [80, 106, 87, 126], and intrinsically gapless topological phases [100, 119].
A crucial open question is how the concept of SEQC manifests in nH systems. Progress on this front remains challenging because studying generic nH many-body systems is notoriously difficult, plagued by numerical instabilities and inaccuracies that obscure the underlying physics [127, 128, 16, 68, 129, 130].
@@
To make substantive progress, we adopt a strategy centered on exactly solvable models. We systematically explore $\mathcal{PT}$-symmetry–enriched non-unitary criticality in a family of one-dimensional nH free-fermion models, uncovering a new class of non-unitary critical points that host robust topological edge modes. Using the combination of exact solution and numerical simulations, we show that although all critical points in these models are described by non-unitary CFTs with central charge c=2c=-2, at least one of them is further enriched by $\mathcal{PT}$ symmetry, giving rise to topologically non-unitary criticality with stable edge states. Remarkably, the associated topological degeneracies are encoded in the purely imaginary part of the entanglement entropy scaling, a feature unique to nH systems. Finally, we elucidate the emergence of topological edge states at this critical point by proposing an entirely new mechanism—generalized mass inversion—that is absent in Hermitian counterparts.
@@
which depends only on $v_{k}$. The vertical dashed line $w=v$ in Fig. 1(a) therefore marks a change in edge-mode count that is independent of $u$. Note that inside the $\mathcal{PT}$-broken wedge the dashed line is not a phase boundary, because the Berry (Zak) phase, which determines the phase structure, no longer corresponds to the winding number there [132, 28, 133].
@@
First we review the ordinary ($\alpha=1$) nH SSH model. At QCP wv=uw-v=-u, the correlation eigenvalues are real and occur in $(\nu,1{-}\nu)$ pairs, ensuring that the entanglement entropy remains real and scales with central charge c=2c=-2, consistent with a non-unitary CFT (see Fig. 2(a)).
In contrast, the QCP wv=uw-v=u exhibits a complex conjugate pair of correlation eigenvalues, $\nu_{\pm}=1/2\pm i\mathcal{I}$, which lead to an imaginary contribution to the entanglement entropy. While previous studies [36] encountered ambiguities due to the multivalued complex logarithm, we restore the correct analytic structure by consistently shifting one logarithmic branch by $2\pi$. This yields an additional term $i\pi$ in the entropy, while preserving the scaling form with c=2c=-2 in the real part (see Fig. 2(b)). Unlike in Hermitian systems, the entropy here acquires a constant imaginary part; we defer its physical interpretation to a later section. Full details of the branch-cut prescription are provided in the Supplementary Material (SM) [133].
@@
Entanglement–edge correspondence and the imaginary entropy—We now focus on the topological aspects of these non-unitary critical points. Specifically, at the critical points of the $\alpha$-nH SSH chain, we observe a correspondence between the physical and entanglement spectra. Under OBC, the system supports localized edge modes with purely imaginary energies $E=\pm iu$. Under PBC, the entanglement spectrum features complex-conjugate eigenvalue pairs of the form $\nu=1/2\pm i\mathcal{I}$, whose number matches that of the OBC edge modes [see Fig. 3(a,b)]. Using the standard correlation–entanglement mapping $\nu=(1+e^{\epsilon})^{-1}$ [36], equivalently
@@
corresponding to a purely imaginary entanglement energy. This generalizes the Li–Haldane picture [134]: although both energy are no longer real or degenerate, the one-to-one counting and the vanishing real parts ($\operatorname{Re}[E]=\operatorname{Re}[\epsilon]=0$) still match.
@@
A key consequence concerns the imaginary part of the bulk entanglement entropy under PBC. Requiring the correct c=2c=-2 scaling at all critical points forces us to adopt the same branch-cut convention used in the previous section. With this convention, each complex-conjugate pair contributes $i\pi$ to $S_{A}$. Together with the entanglement spectrum– edge correspondence, at a QCP with winding number $\omega$,
@@
This identifies a subleading imaginary contribution $i\pi\omega$. Unlike in Hermitian SEQC, where topological degeneracies can be diagnosed via the Affleck–Ludwig boundary $g$ function [80] in the entanglement entropy under OBC and entanglement spectrum under PBC [105, 87, 126], the nH case renders them directly visible in the imaginary part of the bulk entanglement entropy—a feature unique to nH systems.
It is then natural to extend the same branch choice into the adjacent SPT phase by continuity. With that extension, $\operatorname{Im}[S_{A}]$ takes the same constant value throughout each fixed-$\omega$ sector. Thus $\operatorname{Im}[S_{A}]$ serves as a topological index reflecting the winding number $\omega$ across the $\mathcal{PT}$-symmetric sector.
@@
Generalized mass inversion in $\mathcal{PT}$ symmetry-enriched non-Hermitian QCPs—A distinctive feature of the $\alpha$-nH SSH model is the appearance of an intermediate $\mathcal{PT}$-broken symmetry-breaking phase absent in Hermitian free-fermion systems [76]. This additional phase shifts the location of the critical point and is closely tied to the persistence of edge modes across the transition. In the Hermitian $\alpha\geq 2$-chain, topological edge states rely on kinetic inversion [78] with long-range hopping to maintain edge states at criticality. By contrast, in the nH SSH model, the survival of edge modes follows from a simple generalization of mass inversion without requiring long-range hopping ($\alpha=1$).
@@
decouple at $E=\pm iu$, yielding localized edge solutions with decay length $\sim 1/m$ for $m>0$ even as the spectrum gap $\sqrt{m^{2}-u^{2}}$ closes.
@@
More generally, when more than one mass term $m_{a}$ appears in the Hamiltonian, take $M=\sum_{a}m_{a}\sigma_{a}$. There are two scales that control the physics here: (i) the decay rate $\kappa$, which is set by the domain-wall mass—the component that flips sign across the boundary—so $\kappa\simeq-|m_{\mathrm{bdy}}|$; (ii) the spectrum gap $\Delta$, which is set by the total mass, $\Delta=\sqrt{\sum_{a}m_{a}^{2}}$. Thus we can see that these two scales can decouple when there are multiple mass terms. We refer to this scale decoupling, together with ordinary mass inversion at the boundary, as generalized mass inversion.
For a nH SSH model, an imaginary mass $iu$ comes in so that $\Delta=\sqrt{m^{2}-u^{2}}$. Tuning $u\to m$ closes the bulk gap $(\Delta\to 0)$ while keeping the domain-wall mass $m$ finite, hence realizing a SEQC. In contrast, in Hermitian models there is no imaginary mass, so the gap cannot close without delocalizing the edge mode; this route is therefore unique to the non-Hermitian SEQC setting, making a free-fermion SEQC without long-range hopping possible only in the nH case. For higher $\alpha$, multiple coefficients enter the edge equation, and the neat single-parameter inversion picture no longer applies (see the SM [133]).
@@
The schematic in Fig. 4(b) depicts an interface geometry (left: trivial gapped phase; right: nH SSH at its topological QCP). Unlike the OBC case, the interface equations remain coupled; while the lattice calculation is complex, we obtain an explicit continuum (field-theory) solution demonstrating that generalized mass inversion pins an edge state at the interface [133].
@@
Discussion and concluding remarks.—To summarize, we uncover a new class of non-unitary criticality enriched by $\mathcal{PT}$ symmetry, which hosts robust topological edge states. Specifically, we construct a broad family of one-dimensional $\mathcal{PT}$-symmetric nH free-fermion models with $\alpha$-range hopping, all exhibiting critical points described by non-unitary CFTs with central charge c=2c=-2, as confirmed numerically through the scaling of entanglement entropy and ground-state energy. More importantly, we unambiguously demonstrate that these nH critical points support nontrivial edge states protected by $\mathcal{PT}$ symmetry. The associated topological degeneracies ars reflected in the bulk entanglement spectrum—representing a generalized Li-Haldane correspondence at nH criticality—and, remarkably, is also encoded in the imaginary part of the entanglement entropy, a feature unique to nH systems. Finally, we analytically show that the topological edge modes at these QCPs are stabilized by a new mechanism, generalized mass inversion—without requiring long-range hopping—highlighting a distinct protection mechanism from Hermitian counterparts.
Looking ahead, identifying symmetry-enriched nH criticalities in higher dimensions and in interacting many-body systems represents an interesting direction for future work. Regarding experimental realization, non-Hermitian quantum systems can be simulated on photonic platforms, where gain and loss are experimentally implemented [135, 136, 137, 138]. Although the main observables on these platforms typically relate to single-particle physics, the survival of topologically protected edge modes within a fine-tuned gapless spectrum can still be experimentally tested [139]. Furthermore, entanglement properties of non-Hermitian many-body states could potentially be measured in spin systems (via a Jordan-Wigner transformation) using post-selection [140]. This investigation, however, is a direction for future work beyond the scope of our study. Our work not only generalizes the notion of topology to the realm of non-unitary criticality but also contributes to a deeper understanding of critical phenomena in nH systems.